# Navigating the Ethics of Internet Measurement: Researchers’ Perspectives from a Case Study in the EU

### v1

@@ 4.1.3 Difficult to be Transparent and Accountable
P4: “in terms of honey pots the idea is not to be transparent… I have done DPI  [^6] Deep Packet Inspection in the past it’s not very transparent in the first place.”
@@
Researchers mentioned the difficulties in communicating with the operators (P6) who can sometimes get angry and frustrated (P16). In an instance where measurement studies crashed mail servers, operators “became a bit abusive towards the researchers” (P16). Prior work has noted that system operators often work in fast-paced high-pressure environments where they are responsible for uninterrupted system operations, have to quickly resolve security related issues and also support end-users [kaur2022needed]. They are often overworked and overburdened, while at the same time their work remains invisible [Kaur23]. Such work dynamics can lead to frustrations, specially when an external entity starts to interact with your infrastructure without your knowledge.
@@
P2: “If I break some other company’s network, I just have to say, ”apologies. From my deep heart, I apologize, because I can’t do anything.”
@@ 4.1.4 Uncertainty About Ethical Boundaries
P16: “If you didn’t do it and just write some general blah blah blah blah which also a lot of people do and the reviewers notice you usually get guided by the reviewers to write some blah blah blah blah which actually makes sense. So it’s like look we’re doing something but it’s like it doesn’t feel like like deep change within the community recognizing that this is something important they have to do it’s more like a performative action. Like we did the ethics dance if something went wrong it’s not our fault.”
@@
However, ethics and legalities “tend to be two different things” (P16), and legal boundaries are not always unambiguous either. Laws governing Internet measurements were generally not written with benign research scans in mind, and they vary widely across jurisdictions. Existing literature underscores this profound legal uncertainty [Florian]. Some participants were not clear about the legal implications of their work, some delegated the responsibility to the third-party data providers and some assumed that they were complying with legal regulations. However, something permissible in one jurisdiction might be illegal in another, creating a “very dicey” (P8) situation.
@@
Some described operating under the assumption of research exception or academic freedom, but with significant uncertainty about the scope and limitations of such protections. P10, conducting active SSH measurements, explained:
@@
A censorship researcher shares:
P11: “Usually in the country where internet censorship is present, such as China or Russia or Saudi Arabia, it is by law. So, if you’re circumventing this or looking at the censorship, how it works, […] it’s illegal. So, you are technically breaking the law in that country.”
Legitimate research can lie in a legal gray area. The overall responsibility then lies on the researchers who must decide on how to get on with this situation, from where and whom to take advice, and how to publish the findings of their work without putting anyone in legal jeopardy.
@@
Another participant had a similar issue while scanning the SSH off ports that “triggered so many annoyed people” (P10). They further explained the frustration, anger and shock of the concerned party as well:
@@
Researchers “do not have access to a capable IRB” (P16) always. Hence, regardless of ethics approval, one still must be considerate of the potential harms and make the call because “in the end it’s your fault […] it’s always you who does it” (P16). Researchers suggested mixed expert panels or consulting domain experts - similar to prior work [Fiebig2024] - could help both ethicists and researchers make decisions.
@@ 4.1.5 Hurdles in the Ethics Review Process
Ethics review boards often lack domain-specific expertise [Fiebig2024].
@@
Recent work has noted that measuring Internet connectivity might actually bypass government censorship [wendzel2025survey], which could put people in legal trouble or put their safety at risk. Without knowing these hidden details, review boards cannot properly check whether the study is ethically sound or not. Simple technical descriptions may hide deep ethical issues. P16 referenced a study [Encore] that used people’s browsers to check for censorship by trying to access blocked websites and has become a key example in talks about the limits of IRB reviews in computer science research especially in Internet measurement.
@@ 4.2.1 Frameworks and Guidelines
The Menlo Report [Menlo2012] is “rather old” (P10) but is the “baseline thing” (P16).
P10: “And that report tells you, okay, you have to clearly state who you are, what you’re doing. So we had all of this enabled, the reverse DNS. We had a website that if you’re gonna look and see who this IP is, there’s a website explaining what we’re doing and why we are doing.”
@@ 4.2.2 Ethical Reporting in Publications
However, depending on the type of Internet measurement, ethics considerations can change drastically.
P11: “I think censorship is more strict towards ethics. Traffic analysis of course also has an ethics component, but […] for traffic analysis you’re not going to be using a singular user’s traffic pattern. It’s going to be bundled and anonymized by hundreds of millions of users, so really there are not that many privacy concerns left at that point. […] we need actual individual people to help us, so the ethics for censorship I think are a bit more strict than for traffic analysis. This also reflects in my papers, like my traffic papers have a four or five sentence ethics concern section, and my censorship papers have an entire page, so there’s a lot more work involved.”

### v2

@@ 4.1.3 Difficult to be Transparent and Accountable
P4: “I have done DPI [^6] Deep Packet Inspection in the past it’s not very transparent in the first place.”
@@
Researchers mentioned the difficulties in communicating with the operators (P6) who can sometimes get angry and frustrated (P16). In an instance where measurement studies crashed mail servers, operators “became a bit abusive towards the researchers” (P16). Prior work has noted that system operators often work in fast-paced high-pressure environments where they are responsible for uninterrupted system operations, have to quickly resolve security related issues and also support end-users [15]. They are often overworked and overburdened, while at the same time their work remains invisible [16]. Such work dynamics can lead to frustrations, specially when an external entity starts to interact with your infrastructure without your knowledge.
@@
P2: “If I break some other company’s network, I just have to say, “apologies. From my deep heart, I apologize”, because I can’t do anything.”
@@ 4.1.4 Uncertainty About Ethical Boundaries
P16: “If you didn’t do it and just write some general blah blah, which also a lot of people do, and the reviewers notice, you usually get guided by the reviewers to write some blah blah blah which actually makes sense. So it’s like look we’re doing something but it’s like it doesn’t feel like like deep change within the community recognizing that this is something important they have to do. It’s more like a performative action. Like we did the ethics dance if something went wrong it’s not our fault.”
@@
However, ethics and legalities “tend to be two different things” (P16), and legal boundaries are not always unambiguous either. Laws governing Internet measurements were generally not written with benign research scans in mind, and they vary widely across jurisdictions. Existing literature underscores this profound legal uncertainty [12]. Some participants were not clear about the legal implications of their work, some delegated the responsibility to the third-party data providers and some assumed that they were complying with legal regulations. However, something permissible in one jurisdiction might be illegal in another, creating a “very dicey” (P8) situation.
@@
Some described operating under the assumption of research exception or academic freedom, but with significant uncertainty about the scope and limitations of such protections. P10 explained:
@@
One researcher shares that “usually in the country where Internet censorship is present, such as China or Russia or Saudi Arabia, it is by law. So, if you’re circumventing this or looking at the censorship, how it works, […] it’s illegal. So, you are technically breaking the law in that country.” Legitimate research can lie in a legal gray area. The overall responsibility then lies on the researchers who must decide on how to get on with this situation, from where and whom to take advice, and how to publish the findings of their work without putting anyone in legal jeopardy.
@@
Another participant described a similar issue when conducting active network scans on off-ports which “triggered so many annoyed people” (P10). They further explained the frustration, anger and shock of the concerned party as well:
@@
Researchers “do not have access to a capable IRB” (P16) always. Hence, regardless of ethics approval, one still must be considerate of the potential harms and make the call because “in the end it’s your fault […] it’s always you who does it” (P16). Researchers suggested mixed expert panels or consulting domain experts - similar to prior work [10] - could help both ethicists and researchers make decisions.
@@ 4.1.5 Hurdles in the Ethics Review Process
Ethics review boards often lack domain-specific expertise [10].
@@
Recent work has noted that measuring Internet connectivity might actually bypass government censorship [28], which could put people in legal trouble or put their safety at risk. Without knowing these hidden details, review boards cannot properly check whether the study is ethically sound or not. Simple technical descriptions may hide deep ethical issues. P16 referenced a study [5] that used people’s browsers to check for censorship by trying to access blocked websites and has become a key example in talks about the limits of IRB reviews in computer science research especially in Internet measurement.
@@ 4.2.1 Frameworks and Guidelines
The Menlo Report [17] is “rather old” (P10) but is the “baseline thing” (P16). It “tells you, okay, you have to clearly state who you are, what you’re doing” (P10), e.g., having “a website that if you’re gonna look and see who this IP is, there’s a website explaining what we’re doing and why we are doing” (P10).
@@ 4.2.2 Ethical Reporting in Publications
However, depending on the type of Internet measurement, ethics considerations can change drastically. For instance, while traffic analysis “has an ethics component” and papers often include “a four or five sentence ethics concern section”, “censorship [research] is more strict towards ethics” involving “a lot more work” and a more detailed ethics section.